1.A spinner has 4 equal sectors colored red, yellow, blue, and green. What is the probability of landing on a sector that is not red?
Answer #1

P(Red) = P(Yellow) = P(Blue) = P(Green) = 1/4

P(Not red) = 1 - P(Red)

= 1 - 1/4

= 3/4 ( = 0.75)